Circular No. 8592

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005dq, 2005dr, 2005ds Three apparent supernovae have been reported from unfiltered CCD

images — 2005dq by M. Armstrong (cf. IAUC 8406) and the other two by D. R. Madison, R. R. Prasad, and W. Li (LOSS/KAIT; cf. IAUC 8590).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005dq Aug. 30.086 22^{h}44^{m}46.^{s}45 +33^{o}27′37′′.6 17.9 9′′.4 E, 0′′.5 S
2005dr Aug. 29.51 5 02 17.18 + 7 38 21.4 18.5 12′′.0 W, 12′′.0 N
2005ds Aug. 31.21 17 57 39.89 +27 50 18.4 18.7 14′′.4 W, 13′′.1 N

Additional approximate magnitudes from the respective observers: 2005dq in UGC 12177, 1991 Sept. 13, [20.8 (Palomar Sky Survey red plate); 1992 Oct. 19, [22.5 (POSS blue plate); 2003 Sept. 21, [19.5 (Armstrong); 2005 Aug. 30.895 UT, 17.9. 2005dr in UGC 3229, 2004 Mar. 8.18, [19.5; 2005 Aug. 31.49, 18.6. 2005ds in UGC 11064, Aug. 22.21, [19.5; 26.19, [19.0; Sept. 1.20, 18.5.

(76818) 2000 RG_{79}

B. D. Warner, Colorado Springs, CO; P. Pravec, Ondřejov Observatory; and D. Pray, Coventry, RI, report that photometric observations obtained during Aug. 7–28 show that the minor planet (76818) is a binary system with an orbital period of 14.125 ± 0.01 hr. The primary rotates with a period 3.1664 ± 0.0002 hr, and its lightcurve amplitude of 0.14 mag is indicative of a nearly spheroid shape. Mutual eclipse/occultation events that are ∼ 0.14 mag deep indicate a secondary-to-primary mean-diameter ratio of 0.37 ± 0.03.

COMETS C/2005 Q2 AND C/2005 Q3 (SOHO) Following are details on two additional comets found on SOHO C2-

coronagraph website images (cf. IAUC 8585). C/2005 Q2 was a Meyergroup comet, reaching a magnitude of perhaps 6.5 and appearing stellar on all images except two or three, where it appeared elongated. C/2005 Q3 appears to be a non-group comet, peaking at mag perhaps 8.5 when it first entered the field-of-view, but then fading rapidly as it approached the sun.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 Q2 Aug. 23.071 10^{h}02.^{m}6 +13^{o}16′ C2 TH 2005-R06
C/2005 Q3  30.104 10 28.3 +10 21 C2 TH 2005-R06
2005 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green